OpenStudy (anonymous):

Please Help Ill Give Medal and Fan U!!!!!!! The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n)=-250n^2 +2,500n-4,000 Part A: What are the zeros of the above function and what do they represent? Show your work. Part B: Find the maximum profit by completing the square of the function P(n). Show your work. Part C: What is the axis of symmetry of the function P(n)?

OpenStudy (sidsiddhartha):

for part A put P(n)=-250n^2 +2,500n-4,000=0 and u'ii get values of n. i think those will be zeroes of the above function for part B P(n)=-250n^2 +2,500n-4,000 differntiate it p(n)'= -500n + 2500 pur p(n)'=0 u will get n = 5. now If you put 5 into the original function, it becomes -250(5^2) + 2500(5) - 4000 = 2250 for part C i'm assuming u know theory of quadratic equations as the above function is a quadratic equation so it's axis of symmetry=-(b/2*a) here b=2500 and a=-250 so -(b/2*a)=(2500/500)=5 hope u understand